If the 6th term is 18 and the 8th term is 32 what is the 7th term?
Answer by richwmiller(17219)

Which kind of series? A, G or H?
Arithmetic
an = a + (n - 1)d
18 = a + 5d
32 = a + 7d
a = -17, d = 7
a7 = -17 + 6*7
a7=42-17
a7=25
geometric
Tn = t * r^(n - 1)
18 = t * r^(5),
32 = t * r^(7)
r = -4/3, t = -2187/512
and
r = 4/3, t = 2187/512
T = 2187/512 * 4/3^(6)
t=24